OpenStudy (lxelle):

Show that the equation log (x+5) = 2 - log x may be written as a quadratic equation in x.

OpenStudy (anonymous):

\[\log_{10}(x+5)=2-\log(x)=>\log(x+5)+logx=2=>\log(x)(x+5)=2=>x(x+5)=10^2=>x^2+5x=100\]

OpenStudy (lxelle):

I dont understand.

OpenStudy (anonymous):

first bring log(x) other side by adding it on both sides

OpenStudy (anonymous):

then by log (a)+log (b)=log(a.b) property we get log(x)(x+5)

OpenStudy (lxelle):

the answer is supposed to be x2+5x=100

OpenStudy (anonymous):

yes that is the answer (what i typed above is not visible even to me),i wrote a long stright equation

OpenStudy (anonymous):